# Emergent Social Dynamics of LLM Agents in the El Farol Bar Problem

### v1

@@ 3.2 Analysis
To quantitatively verify these qualitative observations, we examined the results of 10 independent simulation runs. We first measured the time difference between clustering and bar crowding across trials (Figure 9). We defined the clustering time ($T_{d}$) as the first moment when over 60% of agents gathered within a certain distance of their group’s average position, and the bar crowding time ($T_{b}$) as when over 60% of agents were inside the bar. Here, this certain distance for calculating $T_{d}$ was set to 10, the same as the side length of the bar. The distribution of the time difference, $\Delta T=T_{b}-T_{d}$, reveals a consistent time lag between these events. The consistent positive distribution indicates that clustering reliably preceded crowding.
Figure 9: Distribution of the time difference, $\Delta T=T_{b}-T_{d}$, over 10 simulations. This histogram shows the time lag between when agents formed a cluster ($T_{d}$) and when the bar became crowded ($T_{b}$). The positive distribution indicates that clustering consistently occurred before crowding.
@@
Next, to understand the decision-making logic behind this equilibrium, we analyzed the distribution of agent actions based on their location (inside/outside the bar) and the bar’s status (crowded/not crowded), as shown in Figure 11. When inside the bar (A), the rate of agents choosing to ‘stay‘ drops significantly from 76.5% to 42.4% when the bar becomes crowded, indicating a clear exit pressure. Conversely, when outside the bar (B), the ‘stay‘ rate dramatically increases from 39.0% to 81.2% when the bar is crowded, suggesting a waiting behavior. This demonstrates that the agents employ context-dependent, rational strategies.
@@ 4 Discussion
Around the 60% threshold, agents repeatedly entered and exited the bar, with attendance fluctuating but stabilizing slightly above the threshold across all trials (Figure 10). This dynamic equilibrium suggests an emergent, autonomous solution to the unstated task of managing congestion. This behavior is driven by distinct rational strategies that differ based on location. As shown in Figure 11, agents outside a crowded bar overwhelmingly choose to stay, effectively waiting for conditions to improve. In stark contrast, agents already inside experience a clear “exit pressure,” significantly increasing their probability of moving. Furthermore, this intent to leave is not random movement; the microscopic analysis in Figure 12 reveals that agents inside a crowded bar consistently exhibit a strong positive direction (away from the bar), which is a goal-oriented behavior toward the exit.
@@ Acknowledgments
Figure 14 illustrates the movement patterns of LLM agents in the library scenario. Unlike the bar scenario where agents frequently exhibited coordinated group formation before entering the target location, the library simulation showed predominantly individual movement patterns. Agents approached the library independently without the spontaneous clustering behavior observed in the bar context (Figure 4, step 250). This suggests that the collective coordination behavior is not merely an emergent property of spatial constraints or communication mechanisms alone, but is influenced by the semantic context of the destination. To understand the fundamental semantic differences driving these behavioral variations, we analyzed the messages generated by agents in both scenarios using word cloud visualizations (Figure 15). The word cloud analysis reveals striking differences in the social language used by agents in each context. In the bar scenario (Figure 15A), the word “Together” appears prominently and frequently in agent communications, indicating a strong emphasis on collective participation and social coordination. Conversely, in the library scenario (Figure 15B), “Together” is notably absent or significantly less prominent, suggesting that agents conceptualize library visits as more individualistic activities.

### v2

@@ 2.1 El Farol Bar Problem
The core mechanic of the simulation revolves around the bar’s comfort level, which is determined by the crowding threshold. If fewer than 60% of all agents are inside the bar, the environment is considered “comfortable.” If the attendance is 60% or more, it becomes “uncomfortable.” Crucially, only agents currently inside the bar receive direct feedback on this status at each step. All agents can communicate with others within their vicinity (a radius of 5, as shown in Table 1), provided they are in the same area (i.e., both inside or both outside the bar).
@@ 3.2 Analysis
To quantitatively verify these qualitative observations, we examined the results of 10 independent simulation runs. We first measured the time difference between clustering and bar crowding across trials (Figure 9). We defined the clustering time ($T_{d}$) as the first moment when over 60% of agents gathered within a certain distance of their group’s average position, and the bar crowding time ($T_{b}$) as when over 60% of agents were inside the bar. Here, this certain distance for calculating $T_{d}$ was set to 10, the same as the side length of the bar. The distribution of the time difference, $\Delta T=T_{b}-T_{d}$, reveals a consistent time lag between these events. The consistent positive distribution indicates that clustering reliably preceded entry into the bar. This social clustering is specific to the bar setting, as shown by the contrasting results from a library scenario (grey bars), which are detailed in Appendix 2.
Figure 9: Distribution of the time difference, $\Delta T=T_{b}-T_{d}$, over 10 simulations. This histogram shows the time lag between when agents formed a cluster ($T_{d}$) and when the bar became crowded ($T_{b}$). The positive distribution for the bar scenario (blue) indicates that clustering consistently occurred before crowding. The results for the comparative library scenario (grey) are discussed in Appendix 2.
@@
Next, to understand the decision-making logic behind this equilibrium, we analyzed the distribution of agent actions based on their location (inside/outside the bar) and the bar’s status (crowded/not crowded), as shown in Figure 11. When inside the bar (A), the rate of agents choosing to “stay” drops significantly from 76.5% to 42.4% when the bar becomes crowded, indicating a clear exit pressure. Conversely, when outside the bar (B), the “stay” rate dramatically increases from 39.0% to 81.2% when the bar is crowded, suggesting a waiting behavior. This demonstrates that the agents employ context-dependent, rational strategies.
@@ 4 Discussion
Around the 60% threshold, agents repeatedly entered and exited the bar, with attendance fluctuating but stabilizing slightly above the threshold across all trials (Figure 10). This dynamic equilibrium suggests an emergent, autonomous solution to the unstated task of managing congestion. This behavior is driven by distinct rational strategies that differ based on location. As shown in Figure 11, agents outside a crowded bar overwhelmingly choose to “stay,” effectively waiting for conditions to improve. In stark contrast, agents already inside experience a clear “exit pressure,” significantly increasing their probability of moving. Furthermore, this intent to leave is not random movement; the microscopic analysis in Figure 12 reveals that agents inside a crowded bar consistently exhibit a strong positive direction (away from the bar), which is a goal-oriented behavior toward the exit.
@@ Acknowledgments
This contextual difference is quantitatively evident in the timing of agent clustering versus entry (Figure 9). In the bar scenario, agents consistently form social clusters before the location becomes crowded ($\Delta T>0$), indicating a preparatory social phase. In stark contrast, the library scenario shows no such coordinated waiting period; agents tend to enter individually without prior clustering, resulting in $\Delta T$ values centered around zero. This confirms that the observed social coordination is not a generic behavior but is triggered by the cultural associations of a “bar” as a social venue.
This quantitative finding is visually supported by the agents’ movement patterns, as illustrated in Figure 14. Unlike the bar scenario where agents frequently exhibited coordinated group formation before entering the target location, the library simulation showed predominantly individual movement patterns. Agents approached the library independently without the spontaneous clustering behavior observed in the bar context (Figure 4, step 250). This suggests that the collective coordination behavior is not merely an emergent property of spatial constraints or communication mechanisms alone, but is influenced by the semantic context of the destination. To understand the fundamental semantic differences driving these behavioral variations, we analyzed the messages generated by agents in both scenarios using word cloud visualizations (Figure 15). The word cloud analysis reveals striking differences in the social language used by agents in each context. In the bar scenario (Figure 15A), the word “Together” appears prominently and frequently in agent communications, indicating a strong emphasis on collective participation and social coordination. Conversely, in the library scenario (Figure 15B), “Together” is notably absent or significantly less prominent, suggesting that agents conceptualize library visits as more individualistic activities.